OpenStudy (anonymous):

I am having so much trouble with my calculus homework please someone help me

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

the1st question is Determine the absolute value of each complex number |-3i|

OpenStudy (anonymous):

okay, so i = square root of -1 right

OpenStudy (anonymous):

and the absolute value of -3 is 3

OpenStudy (anonymous):

so it's probably 3i .... i think that's what they're asking.

OpenStudy (anonymous):

this is so hard i don;t remember doing this in high school the problem i'm having is the showing of the work i have absolutly no idea what ia what

OpenStudy (anonymous):

Don't worry. You just need to follow the examples in the book and try doing them yourself. If you still don't get it, go to khanacademy.org.. he's really good =) And in this case, you don't need to show that much work except simplify.

OpenStudy (anonymous):

i think you should treat it as finding distance of complex number. |0-3i|= sqrt (0^2 + (-3)^2) = sqrt (9) = 3 :)

OpenStudy (anonymous):

Yeah, you're right melor.

OpenStudy (anonymous):

that looks like alien writing i'm so confused

OpenStudy (anonymous):

you may refer here http://www.clarku.edu/~djoyce/complex/abs.html

OpenStudy (anonymous):

i'll try that website. Thanks!!!!! I feel so dumb this is the only class that i'm failing

OpenStudy (anonymous):

don't worry, maths is about experiencing the problems. just keep trying and one day you'll get over it ;)

OpenStudy (anonymous):

thanks!!